# An adaptive finite element multigrid solver using GPU acceleration

### v1

@@ 1 Introduction
These challenges are one of the central reasons that parallel accelerator co-processors, such as GPU’s and TPU’s, are still only rarely used in numerical simulations, including for adaptive multi-grid methods. Accelerators are highly attractive since the computational power of CPUs has stagnated in the last decade while those accelerators has increased dramatically.
Accelerators use task- and data-parallelism as well as specialization to achieve a very large peak performance. While a high degree of parallelism is inherent in some problems, e.g. in computer graphics and deep learning, it poses challenges for the implementation of adaptive algorithms in numerical linear algebra and scientific computing. This holds true in particular for data-parallelism that requires regular, structured computations with uniform memory access patterns. One direction to address this gap is by changing to algorithms with highly regular computations, such as lattice Boltzmann methods in fluid mechanics [25, 27]. A minimally invasive use of GPU’s in classical finite element simulation tools is also possible but yields little to no acceleration [14].
Another challenge for the the use of accelerator hardware in scientific computing is that these are optimized for single, half or even less precision while double precision remains the standard for most calculations in numerical mathematics. In [29], this is addressed by using a hierarchical approach to modify the linear systems, denoted as pre-handling. An alternative direction is the use of machine learning methods, especially deep neural networks, in the context of simulations, which can be executed very efficiently and in lower precison on accelerator hardware with speedups up to several orders of magnitude, e.g. [7].
@@
The integration of GPU’s into an existing geometric multigrid solver was already presented in [15]. The obtained speedup were moderate but the work describes how a transparent implementation into an existing software can be achieved. The consequences of mixed precision arithmetic were also discussed in this work. In [13], the same authors presented a geometric multigrid method that is based entirely on sparse matrix-vector multiplications and can thus be easily and efficiently implemented on different hardware using suitable libraries. Applied to stationary linear differential equations, an acceleration of the overall solver by a factor of up to 5 was achieved compared to CPU systems. In [17], the authors investigate the GPU parallelization of several smoothers within a matrix-based geometric multigrid method. The implementation allows for locally refined unstructured meshes, but, the obtained speedup is limited and does not substantially exceed that of multi-core CPU’s. NVIDIA [32] has presented a highly efficient geometric multigrid solver for 3d linear elliptic problems. It reaches very high throughput and demonstrates the potential of native CUDA implementations. It is, however, limited in terms of flexibility and its applicability to unstructured meshes.
@@ 2.1 Finite element discretization
We assume shape regularity in the sense that $\|\nabla T_{T}\|\cdot\|\nabla T_{T}^{-1}\|\leq c$ uniform in $h>0$, see [28, Sec. 4.2]. Structural regularity is relaxed by allowing local mesh refinement with at most 1 hanging node per face, again, see [28, Sec 4.2] for the specific realization in the finite element library Gascoigne 3d [10].
Systems or partial differential equations such as the Navier-Stokes equations or elasticity will be discretized with equal-order finite elements, i.e. $\mathbf{V}_{h}=[V_{h}^{(r)}]^{c}$, where $c\in\mathbb{N}$ is the number of components. Gascoigne 3d combines these components locally. Taking the 3d Navier-Stokes equations as example this means that the vector is represented as a matrix with entries $x_{i,c}$ where $i$ refers to the grid node and $c$ to the component ($c=1$ pressure, $c=2,3,4$ velocities). The system matrix also has double indexing. Each matrix entry $A_{ij}$ is itself a (dense) matrix $A_{ij}\in\mathbb{R}^{n_{c}\times n_{c}}$ with $n_{c}=4$ in the case of 3d Navier-Stokes. Basis functions are always scalar, such that a Navier-Stokes solution is given as
$\begin{pmatrix}p_{h}(x)\\ \boldsymbol{u}_{h}(x)\end{pmatrix}=\sum_{i=1}^{N}\begin{pmatrix}x_{i,1}\\ x_{i,2}\\ x_{i,3}\\ x_{i,4}\end{pmatrix}\phi_{h}^{i}(x).$
Keeping the solution components together is beneficial in terms of cache efficiency for problems where one obtains large system matrices like Navier-Stokes ($c=4$) or 3d elasticity, where $c=6$ when the velocity and deformation fields are combined [8]. The memory layout of this approach is not easily transferred to GPUs as standard libraries like cuSPARSE do not support such blocked matrices. Section 4.1.1 will give details.
@@ 2.2 Adaptivity and geometric multigrid hierarchy
Adaptivity of the computations is realized by means of hierarchical local mesh refinement. If an element $T\in\Omega_{h}$ is chosen for refinement, it is replaced by 4 elements in 2d and 8 elements in 3d. We allow a level jump of 1 for neighboring elements. Refinement of one element can therefore lead to adjacent further refinements. On edges and faces with a level jump, interior degrees of freedom, so-called hanging nodes, are replaced by the interpolation from the neighboring indices. These interpolations are realized by matrix-vector products with sparse matrices that depend on the mesh structure and the chosen finite element degree. Mesh refinement can be based either on a priori knowledge or on a posteriori error estimates, the latter usually in the context of the dual weighted residual method [6, 9].
@@
Successive mesh refinement results in a hierarchy of meshes. Instead of using this hierarchy for the setup of the geometric multigrid solver, we start with the finest mesh $\Omega_{h}=:\Omega_{L}$ and recursively coarsen it until we reach a coarse mesh $\Omega_{H}=:\Omega_{0}$ with small complexity $N_{H}$. In each coarsening step, as many refinements as possible are taken back, yielding a hierarchy $\Omega_{h}=\Omega_{L}\succeq\Omega_{L-1}\succeq\dots\Omega_{0}=\Omega_{H}$. We follow a global multigrid approach where each intermediate level $\Omega_{l}$ spans the complete domain, see [4]. Fig. 1 provides a sketch of the adaptive mesh refinement and the global coarsening procedure that result in two different hierarchies of meshes.
@@ 3 Data-parallel accelerators
Accelerator co-processors, such as GPUs from NVIDIA and AMD and TPUs from Google, are used in a wide range of applications, such as computer gaming and neural networks training. Due to their very large compute power, they also play an increasingly important role in scientific computing. The large compute power is achieved through a very high degree of parallelization as well as specialization. For example, NVIDIA’s GPUs combine MIMD parallelism with up to 128 streaming multi-processors with data-parallelism with a logical, pipelined width of $1024$ and a hardware width of $32$ (and $8$ on the latest hardware). Additionally, significant compute power is provided by dedicated matrix-matrix multiplications engines known as tensor cores and introduced for neural network training. Accelerator co-processors also differ from conventional processors through native hardware support of lower precision data types such as half precision and even 8-bit computations on the latest generation of chips. Accelerator co-processesors as used in this work come with their own, dedicated memory hierarchy starting from a RAM, and with typically two layers of caches. To perform computations on the accelerator, data hence needs to first be transferred to the accelerator RAM through a comparatively slow memory interface.
@@
For specific applications, such as linear algebra or discrete Fourier transform, also highly optimized libraries exist. These provide very high performance while, through the targeted use case, also allow for much simpler usage than, e.g., CUDA. In our work, we will build on a linear algebra formulation of our adaptive PDE solvers. This allows us to use the cuBLAS [23] and cuSPARSE libraries [24] that provides high performance while being simple to use with a software interface that is roughly comparable to those of the conventional BLAS library. Furthermore, cuBLAS and cuSPARSE can still be combined with native CUDA code when necessary, i.e. when an operation cannot be expressed efficiently in linear algebra.
@@ 4 CUDA multigrid
In this section we discuss our extension of the adaptive finite element library Gascoigne 3d to use GPU accelerators for a wide range of computations. The extensions have been designed with two use cases in mind: first, non-stationary but linear systems of partial differential equations, where a single system matrix is reused in every time step or at least in many time steps, and, second, nonlinear problems that can be formulated in terms of fixed matrices that do not need re-assembly. The second cases includes a pressure-projection based solvers for the nonstationary nonlinear Navier-Stokes equations but can be extended to more complex flow problems such as the Boussinesq approximation [12]. Both problem classes have in common that matrices are assembled only once and the complete workflow can be formulated in terms of matrix vector products and the solution of linear systems of equations. Algorithms 2 and 3 show the typical workflow for the problem types. Computations highlighted in blue are completely performed on the GPU, orange marks transfer of data between GPU and CPU.
@@ 4.1.1 Storage
Gascoigne 3d uses a block-wise memory concept in which the various components of a partial differential equation (e.g. the 3 deformation unknowns in 3d solid mechanics or the pressure and the three velocity components in fluid mechanics) are stored next to each other. Matrices are double-indexed and the outer index $(i,j)$ refers to the mesh node yielding the entry $a_{i,j}\in\mathbb{R}^{c\times c}$ that stores (aligned in memory) the local couplings between the solution components as a matrix. In general, this setup helps to efficiently use caching [8, 20]. Using cuSPARSE this block-matrix format is not supported. While vectors are directly transferred to the GPU using the same memory layout, the inner matrix blocks must be resolved and matrices are stored in the usual CSR format. Fig. 2 shows a visualization of the matrix storage concept on the CPU (left) and the GPU (right).
Figure 2: Storage of matrix on CPU (left) and GPU (right). While we employ a block-wise ordering of the matrix entries on the CPU and cluster the $c\times c$ components of a system of PDEs on the CPU, the entries are stored in standard CSR format on the GPU.
@@ 4.1.3 Custom CUDA kernels
The flexibility to combine cuBLAS, cuSPARSE an native CUDA is, in our opinion, an important feature to avoid CPU-GPU data transfer and to be able to flexibly implement algorithms with CUDA-Gascoigne-3d. The total number of required custom kernels is very small and serves as guideline for measuring the effort to port further applications to the GPU.
@@ 4.1.4 Geometric Multigrid CUDA
As a concrete example, we describe in the following how geometric multi-grid algorithm in Algo. 1 is realized in CUDA-Gascoigne-3d. All computations are entirely performed on the GPU. Hence, data transfer overhead is incurred only at the beginning and end of the computations, see also Algorithm 2 and 3.
@@
with $P_{l}\in\mathbb{R}^{N_{l+1}\times N_{l}}$ and its entries given by $\chi_{ij}$. With the number of distinct elements in $P_{l}$ being very small, the use of the matrix formulation would be sub-optimal on the CPU. On the GPU, however, it is performant and allows us to use cuSPARSE. The multigrid restriction $R_{l}:\Omega_{l+1}\to\Omega_{l}$ is the transpose of the prolongation, i.e. $R_{l}=P_{l}^{T}$ and hence also implemented using a sparse matrix in cuSPARSE. Similarly to the mesh transfer, also for hanging nodes we formulate the operations for averaging and distributing the values of the solution vector as sparse matrix-vector products, compare Section 2.2.
@@ 4.1.5 GMRES as linear solver on the GPU
For reasons of numerical stability, e.g. on non-uniform meshes, the multigrid iteration can often not be used directly as linear solver. They are then employed as preconditioner in a GMRES iteration. Our GMRES solver follows the approach described in [30, Section 6.5.3], using the modified Gram Schmidt algorithm for orthogonalization and Givens rotation for solving the resulting overdetermined linear system. Due to the multigrid preconditioning, we avoid stability issues in the Gram Schmidt orthogonalization and never have to perform more than 5-10 GMRES steps. The Gram Schmidt iteration is also completely run on the GPU using cuSPARSE and cuBLAS operations. As a maximum of $n_{G}\leq 10$ steps is usually performed, the resulting overdetermined system used to find the GMRES solution is very small and its step-wise transformation to a diagonal matrix with Givens rotations is performed on the CPU. The size of this matrix does not exceed $\mathbb{R}^{(n_{G}+1)\times n_{G}}$ such that memory transfer between GPU and CPU is negligible. Details are given in [30].
@@ 5 Results
As a third example, we consider the Navier-Stokes equations. For these, we first derive an explicit pressure-correction method, which can be represented by matrix-vector multiplications and a pressure-Poisson problem, latter one being solved with the multigrid solver. Special attention is given to the nonlinearity of the Navier-Stokes equations. Through a reformulation, also this will be approximated by a product with a pre-computed sparse matrix and can easily be performed on the GPU.
If not noted otherwise, all CPU computations are carried out on an Intel Xeon E5-2640 running at 2.40 Ghz using 8 threads. The GPU computations are performed on the same machine using an Nvidia V100 card with 16 GB of GPU RAM. When using the GPU to accelerate the multigrid solver, those parts of Gascoigne that reside on the CPU still run on 8 parallel threads. We always indicated wall-clock times and separately specify the different contributions of Algorithms 2 and 3.

### v2

@@ 1 Introduction
These challenges are one of the central reasons that parallel accelerator co-processors, such as GPUs and TPUs, are still only rarely used in numerical simulations, including for adaptive multi-grid methods. Accelerators are highly attractive since the computational power of CPUs has stagnated in the last decade while those accelerators has increased dramatically.
Accelerators use task- and data-parallelism as well as specialization to achieve a very large peak performance. While a high degree of parallelism is inherent in some problems, e.g. in computer graphics and deep learning, it poses challenges for the implementation of adaptive algorithms in numerical linear algebra and scientific computing. One direction to address this gap is by changing to algorithms with highly regular computations, such as lattice Boltzmann methods in fluid mechanics [30, 32]. A minimally invasive use of GPUs in classical finite element simulation tools is also possible but yields little to no acceleration [16].
Another challenge for the the use of accelerator hardware in scientific computing is that these are optimized for single, half or even less precision while double precision remains the standard for most calculations in numerical mathematics. In [34], this is addressed by using a hierarchical approach to modify the linear systems, denoted as pre-handling. An alternative direction is the use of machine learning methods, especially deep neural networks, in the context of simulations, which can be executed very efficiently and in lower precision on accelerator hardware with speedups up to several orders of magnitude, e.g. [7].
@@
The integration of GPUs into an existing geometric multigrid solver was already presented in [17]. The obtained speedup was moderate but the work describes how a transparent implementation into an existing software can be achieved. The consequences of mixed precision arithmetic were also discussed in this work. In [15], the same authors presented a geometric multigrid method that is based entirely on sparse matrix-vector multiplications and can thus be easily and efficiently implemented on different hardware using suitable libraries. Applied to stationary linear differential equations, the solver exhibited up to a fivefold speedup relative to CPU-based systems.
In [19], the authors investigate the GPU parallelization of several smoothers within a matrix-based geometric multigrid method. The implementation allows for locally refined unstructured meshes, but, the obtained speedup is limited and does not substantially exceed that of multi-core CPU’s. NVIDIA [37] has presented a highly efficient geometric multigrid solver for 3d linear elliptic problems. It reaches very high throughput and demonstrates the potential of native CUDA implementations. It is, however, limited in terms of flexibility and its applicability to unstructured meshes.
@@ 2.1 Finite element discretization
We assume shape regularity in the sense that $\|\nabla T_{T}\|\cdot\|\nabla T_{T}^{-1}\|\leq c$ uniform in $h>0$, see [33, Sec. 4.2]. Structural regularity is relaxed by allowing local mesh refinement with at most 1 hanging node per face, again, see [33, Sec 4.2] for the specific realization in the finite element library Gascoigne 3d [10].
Systems of partial differential equations such as the Navier-Stokes equations or elasticity will be discretized with equal-order finite elements, i.e. $\mathbf{V}_{h}=[V_{h}^{(r)}]^{c}$, where $c\in\mathbb{N}$ is the number of components. Gascoigne 3d combines these components locally. Taking the 3d Navier-Stokes equations as example this means that the vector is represented as a matrix with entries $x_{i,c}$ where $i$ refers to the grid node and $c$ to the component ($c=1$ pressure, $c=2,3,4$ velocities). The system matrix also has double indexing. Each matrix entry $A_{ij}$ is itself a (dense) matrix $A_{ij}\in\mathbb{R}^{n_{c}\times n_{c}}$ with $n_{c}=4$ in the case of 3d Navier-Stokes. Likewise, each vector entry $v_{i}$ is itself a vector $v_{i}\in\mathbb{R}^{n_{c}}$. Keeping the solution components together is beneficial in terms of cache efficiency for problems where one obtains large system matrices like Navier-Stokes ($n_{c}=4$) or 3d elasticity, where $n_{c}=6$ when the velocity and deformation fields are combined [8, 23]. The memory layout of this approach is not easily transferred to GPUs as standard libraries like cuSPARSE do not support such blocked matrices. Section 4.1.1 will give details.
@@ 2.2 Adaptivity and geometric multigrid hierarchy
Adaptivity of the computations is realized by means of hierarchical local mesh refinement. If an element $T\in\Omega_{h}$ is chosen for refinement, it is replaced by 4 elements in 2d and 8 elements in 3d. We allow a level jump of 1 for neighboring elements. The actual refinement can therefore extend further into the domain. Faces and edges on elements with a level jump therefore have inner nodes that are unknowns only on one side. These nodes are called hanging nodes and are replaced by interpolations of their direct neighbours in the finite element approach. In the original Gascoigne, hanging node data used for this interpolation is stored in a sparse structure only covering exactly the hanging nodes. As preparation for GPU parallelization, the data is now represented in a sparse matrix. This matrix $\mathbf{H}_{h}$ is the diagonal identity in all rows belonging to standard degrees of freedom, whereas in rows belonging to hanging nodes, the matrix contains the interpolation weights such that the matrix vector product $\mathbf{y}=\mathbf{H}_{h}\mathbf{x}$ performs all interpolations at once. The computational effort is $O(N)$ ($N$ being the number of unknowns), regardless if hanging nodes are present. However, the constant is small and the overhead in relation to a treatment of the hanging nodes only is negligible. The transpose $\mathbf{H}_{h}^{T}$ plays an important role in the assembly of the matrix and variational residuals. If, for example, an integral $f_{i}=(f,\phi_{i})$ is calculated for $i=1,\dots,N$, all test functions $\phi_{i}$, including those in hanging nodes, are processed first. The result vector $\mathbf{f}_{h}=(f_{i})_{i}$ is then multiplied by $\mathbf{H}_{h}^{T}$ so that the test functions belonging to hanging nodes are correctly taken into account.
Mesh refinement can be based either on a priori knowledge or on a posteriori error estimates, the latter usually in the context of the dual weighted residual method [6, 8].
@@
Successive mesh refinement results in a hierarchy of meshes. Instead of using this hierarchy for the setup of the geometric multigrid solver, we start with the finest mesh $\Omega_{h}=:\Omega_{L}$ and recursively coarsen it until we reach a coarse mesh $\Omega_{H}=:\Omega_{0}$ with small complexity $N_{H}$. The mesh $\Omega_{H}$ is either the coarse starting mesh or any finer mesh chosen to give an optimal balance between recursive multigrid performance and fast direct solution of a small problem. By $L$ we denote the resulting number of multigrid layers. In each coarsening step, as many refinements as possible are taken back, yielding a hierarchy $\Omega_{h}=\Omega_{L}\succeq\Omega_{L-1}\succeq\dots\Omega_{0}=\Omega_{H}$. We follow a global multigrid approach where each intermediate level $\Omega_{l}$ spans the complete domain, see [4]. The main advantage of the global multi-grid is its simplicity. An approximation of the entire problem can be created on each grid level. This allows, for example, the introduction of global constraints, e.g. a normalisation of the pressure, $\int\Omega p\,\text{d}x=0$, in the context of incompressible flows or compliance with material balances $\sum_{i}c_{i}=1$ for chemical reactions, see [5]. The disadvantage of the global approach is the greater effort required, as the multigrid smoother covers the entire grid on each level. However, if one analyses the effort in terms of the accuracy that can be achieved with adaptive grids and not in terms of the number of degrees of freedom, it becomes clear that the global approach has no significant disadvantage in practical applications [4, 9, 8].
Fig. 1 provides a sketch of the adaptive mesh refinement and the global coarsening procedure that result in two different hierarchies of meshes.
@@ 3 Data-parallel accelerators
Accelerator co-processors, such as GPUs from NVIDIA and AMD and TPUs from Google, are used in a wide range of applications, such as computer gaming and neural networks training. Due to their very large compute power, they also play an increasingly important role in scientific computing. The large compute power is achieved through a very high degree of parallelization as well as specialization. For example, NVIDIA’s GPUs combine MIMD parallelism with up to 132 streaming multi-processors with data-parallelism with a logical, pipelined width of $1024$ and a hardware width of $32$ (and $8$ on the latest hardware). Additionally, significant compute power is provided by dedicated matrix-matrix multiplications engines known as tensor cores and introduced for neural network training. Accelerator co-processors also differ from conventional processors through native hardware support of lower precision data types such as half precision and even 8-bit computations on the latest generation of chips. Accelerator co-processors as used in this work come with their own, dedicated memory hierarchy starting from a RAM, and with typically two layers of caches. To perform computations on the accelerator, data hence needs to first be transferred to the accelerator RAM through a comparatively slow memory interface.
@@
For specific applications, such as linear algebra or discrete Fourier transform, also highly optimized libraries exist. These provide very high performance while, through the targeted use case, also allow for much simpler usage than, e.g., CUDA. In our work, we will build on a linear algebra formulation of our adaptive PDE solvers. This allows us to use the cuBLAS [28] and cuSPARSE libraries [29] that provide high performance while being simple to use with a software interface that is roughly comparable to those of the conventional BLAS library. Furthermore, cuBLAS and cuSPARSE can still be combined with native CUDA code when necessary, i.e. when an operation cannot be expressed efficiently in linear algebra.
@@ 4 CUDA multigrid
In this section we discuss our extension of the adaptive finite element library Gascoigne 3d to use GPU accelerators for a wide range of computations. The extensions have been designed with two use cases in mind: first, non-stationary but linear systems of partial differential equations, where a single system matrix is reused in every time step or at least in many time steps, and, second, nonlinear problems that can be formulated in terms of fixed matrices that do not need re-assembly. The second case includes a pressure-projection based solvers for the nonstationary nonlinear Navier-Stokes equations but can be extended to more complex flow problems such as the Boussinesq approximation [12]. Both problem classes have in common that matrices are assembled only once and the complete workflow can be formulated in terms of matrix vector products and the solution of linear systems of equations. Algorithms 2 and 3 show the typical workflow for the problem types. Computations highlighted in blue are completely performed on the GPU, orange marks transfer of data between GPU and CPU.
@@ 4.1.1 Storage
Gascoigne 3d uses a block-wise memory concept in which the various components of a partial differential equation (e.g. the 3 deformation unknowns in 3d solid mechanics or the pressure and the three velocity components in fluid mechanics) are stored next to each other. Matrices are double-indexed and the outer index $(i,j)$ refers to the mesh node yielding the entry $a_{i,j}\in\mathbb{R}^{n_{c}\times n_{c}}$ that stores (aligned in memory) the local couplings between the solution components as a matrix. In general, this setup helps to efficiently use caching [8, 23]. Using cuSPARSE this block-matrix format is not supported. While vectors are directly transferred to the GPU using the same memory layout, the inner matrix blocks must be resolved and matrices are stored in the usual CSR format. Fig. 2 shows a visualization of the matrix storage concept on the CPU (left) and the GPU (right).
Figure 2: Storage of matrix on CPU (left) and GPU (right). On the CPU we employ a block-wise ordering of the matrix entries clustering the $n_{c}\times n_{c}$ components of a system of PDEs. On the GPU standard CSR format is used.
@@ 4.1.3 Custom CUDA kernels
The flexibility to combine cuBLAS, cuSPARSE and native CUDA is, in our opinion, an important feature to avoid CPU-GPU data transfer and to be able to flexibly implement algorithms with CUDA-Gascoigne-3d. The total number of required custom kernels is very small and serves as guideline for measuring the effort to port further applications to the GPU.
@@ 4.1.4 Geometric Multigrid CUDA
As a concrete example, we describe in the following how geometric multi-grid algorithm in Algo. 1 is realized in CUDA-Gascoigne-3d. All computations are entirely performed on the GPU. Hence, data transfer overhead is incurred only at the beginning and end of the computations, see also Algorithms 2 and 3.
@@
with $P_{l}\in\mathbb{R}^{N_{l+1}\times N_{l}}$ and its entries given by $\chi_{ij}$. With the number of distinct elements in $P_{l}$ being very small, the use of the matrix formulation would be sub-optimal on the CPU. On the GPU, however, it is performant and allows us to use cuSPARSE. The multigrid restriction $R_{l}:\Omega_{l+1}\to\Omega_{l}$ is the transpose of the prolongation, i.e. $R_{l}=P_{l}^{T}$ and hence also implemented using a sparse matrix in cuSPARSE. Similarly to the mesh transfer, also for hanging nodes we formulate the operations for averaging and distributing the values of the solution vector as sparse matrix-vector products, see Section 2.2.
@@ 4.1.5 GMRES as linear solver on the GPU
For reasons of numerical stability, e.g. on non-uniform meshes, the multigrid iteration can often not be used directly as linear solver. They are then employed as preconditioner in a GMRES iteration. Our GMRES solver follows the approach described in [35, Section 6.5.3], using the modified Gram Schmidt algorithm for orthogonalization and Givens rotation for solving the resulting overdetermined linear system. Due to the multigrid preconditioning, we avoid stability issues in the Gram Schmidt orthogonalization and never have to perform more than 5-10 GMRES steps. The Gram Schmidt iteration is also completely run on the GPU using cuSPARSE and cuBLAS operations. As we use a highly efficient geometric multigrid solver as preconditioner, a maximum of $n_{G}\leq 10$ GMRES steps is usually required. Hence, the resulting overdetermined system for finding the GMRES solution is very small and its step-wise transformation to a diagonal matrix with Givens rotations is performed on the CPU. The cost for transferring these $n_{G}^{2}$ entries between GPU and CPU is small.. Details on GMRES are given in [35].
@@ 5 Results
As a third example, we consider the Navier-Stokes equations. For these, we first derive an explicit pressure-correction method, which can be represented by matrix-vector multiplications and a pressure-Poisson problem, the latter one being solved with the multigrid solver. Special attention is given to the nonlinearity of the Navier-Stokes equations. Through a reformulation, also this will be approximated by a product with a pre-computed sparse matrix and can easily be performed on the GPU.
Figure 3: Strong scaling of the time-stepping parts of the largest problems described in Sections 5.1, 5.2 and 5.3, with up to $2,000,000$ degrees of freedom on an AMD EPYC 7773X CPU.
All CPU computations are carried out on an AMD EPYC 7773X. For the relatively small problem sizes we consider, our scaling test in Fig. 3 shows that there is only a minor benefit from using a large number of threads. The calculations are dominated by sparse matrix-vector multiplications and the efficiency is therefore limited by the memory bandwidth. This has already been shown in a detailed parallel analysis of fluid-structure interactions based on Gascoigne [14]. A transition to matrix-free finite elements—see, e.g., [24]—would thus offer the greatest potential for improving parallel efficiency. However, this would require a fundamental redesign of the current implementation. We therefore limit the OpenMP based parallel code to 8 threads for the comparisons. The GPU computations are performed on the same machine using an NVIDIA H100 PCIe card with 80 GB of VRAM. When using the GPU to accelerate the multigrid solver, those parts of Gascoigne that reside on the CPU still run on 8 parallel threads. We always indicate wall-clock times and separately specify the different contributions of Algorithms 2 and 3.